Buffer concepts for discussion a. In your own words, explain the purpose of an acid-base buffer. b. How do the acid and base parts of a buffer function to support the buffer? c. Describe the common ion effect and how it is related to LeChateliers Principle. Working with Aqueous Solutions of HF and NaF a. What is the pH of a 0.35 M hydrofluoric acid (HF) solution? Ka 6.8x 104 [1.81] b. What is the pH of a solution containing 0.35 M HF and 0.28 M NaF? [3.07]
Answer #1

Part-a: A buffer is a chemical mixture that resists any small change in the pH of the solution and helps in providing a constPart-b: Mechanism of action of buffer solution in resisting pH change of a solution is referred as buffer action. Consider acNaOH added HCI added Na+ + OH- H+ +C1 Buffer solution CH3COOH CH3COO + H+ CH3COONa — CH3C00- + Na CH3COO + Nat Quenching CH3CNow, according to Le Chaterliers principle, the right side of equilibrium is disturbed due to excess SO42-ion. However, in o6.8 * 104 = [x] [x 0.35 x2 = (6.8 x 104) (0.35) = 2.38 x 104 x = 1.54 x 10-2 = [H*] By definition, pH = -log [H*] pH = -log ((0.28) pH = -log (6.8 x 104) + log - [0.35] = -log 6.8 + 4 log10 + log 0.8 = -0.833 +4 -0.097 pH = 3.07